A piece of copper block of mass 24g at 230 oC is placed in a copper calorimeter of mass 60 oC containing 54g of water at 31 oC .assuming heat losses are negligible,calculate the final steady temperature of mixture.( specific heat capacity of water =4200 JKg^1K^1 specific heat capacity of copper =400JKg^1K^1).


1
Expert's answer
2021-02-09T10:41:49-0500

"m_cc_c(T_c-T)=M_cc_c(T-T_w)+m_wc_w(T-T_w)"


"0.024\cdot 400\cdot(503-T)="


"=0.06\cdot 400\cdot(T-304)+0.054\cdot 4200\cdot (T-304)\to"


"T\approx311\ K=38°" . Answer